HAVING A PROBLEM WITH THIS SYSTEM OF EQUATIONS. NOT SURE I HAVE DONE IT RIGHT.
X+2Y+Z=2
5X=Z
Y=-2X

Answer by mananth(16946)

X+2Y+Z=2 ----1
5X=Z --------2
Y=-2X---------3
substitute the value of z in eq 1
x+2y+5x=2
6x+2y=2 ----a
y+2x=0 ---------3
multiply eq. 3 by 2
2y+4x=0 ---------b
solve equation a & b for x subtract b from a
2x=2
x=1
apply for y and z